OpenStudy (mach_15):

3 - √11 over 2 + √11 Write the expression in simplified radical form. Show your work

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[\frac{3-\sqrt{11}}{2+\sqrt{11}}\]

OpenStudy (misty1212):

multiply top and bottom by \(2-\sqrt{11}\)

OpenStudy (tkhunny):

For a problem that says "show your work":, you probably should show some. BTW - careful use of parentheses might make things easier. (3 - √11)/(2 + √11) -- for example

OpenStudy (mach_15):

@misty1212 yes that's what it is

OpenStudy (tkhunny):

See, that sort of confirmation is unnecessary if you go to the little extra effort to use the notation carefully.